### v1

# MambaVision: A Hybrid Mamba-Transformer Vision Backbone

Abstract
We propose a novel hybrid Mamba-Transformer backbone, denoted as MambaVision, which is specifically tailored for vision applications. Our core contribution includes redesigning the Mamba formulation to enhance its capability for efficient modeling of visual features. In addition, we conduct a comprehensive ablation study on the feasibility of integrating Vision Transformers (ViT) with Mamba. Our results demonstrate that equipping the Mamba architecture with several self-attention blocks at the final layers greatly improves the modeling capacity to capture long-range spatial dependencies. Based on our findings, we introduce a family of MambaVision models with a hierarchical architecture to meet various design criteria. For Image classification on ImageNet-1K dataset, MambaVision model variants achieve a new State-of-the-Art (SOTA) performance in terms of Top-1 accuracy and image throughput. In downstream tasks such as object detection, instance segmentation and semantic segmentation on MS COCO and ADE20K datasets, MambaVision outperforms comparably-sized backbones and demonstrates more favorable performance. Code: https://github.com/NVlabs/MambaVision.
1 Introduction
[…]
Figure 1: Top-1 accuracy vs. image throughput on ImageNet-1K. All measurements conducted on A100 GPU with batch size of 128. MambaVision achieves a new SOTA Pareto front.
During the recent years, Transformers [1] have become the de facto architecture in different domains including in computer vision, natural language processing, speech processing and robotics. In addition, the versatility of the Transformer architecture, primarily due to its attention mechanism, and its flexibility makes it highly suitable for multimodal learning tasks, where integrating and processing information from various modalities is essential. Despite these benefits, the quadratic complexity of the attention mechanism with respect to sequence length make Transformers computationally expensive to train and deploy. Recently, Mamba [2] proposed a new State Space Model (SSM) that achieves linear time complexity and outperforms or matches Transformers in different language modeling tasks [2]. The core contribution of Mamba is a novel selection mechanism that enables efficient input-dependent processing of long sequences with hardware-aware considerations.
Recently, a number of Mamba-based backbones [3, 4] have also been proposed to leverage the strengths of its SSM formulation in vision tasks such as image classification and semantic segmentation. However, the Mamba’s autoregressive formulation, while effective for tasks requiring sequential data processing, faces limitations in computer vision tasks that benefit from a full receptive field: (1) Unlike sequences where order matters, image pixels do not have a sequential dependency in the same way. Instead, spatial relationships are often local and need to be considered in a more parallel and integrated manner. Hence, this results in inefficiency for processing spatial data (2) an autoregressive model like Mamba processes data step-by-step, limiting its ability to capture and utilize global context in one forward pass. In contrast, vision tasks often require understanding the global context to make accurate predictions about local regions.
Vision Mamba (Vim) [3] and others have proposed modifications such as bidirectional SSMs to address lack of global context and spatial understanding. While bidirectional SSMs have the potential to capture more comprehensive context, they introduce significant latency due to the need to process the entire sequence before making predictions. Additionally, the increased complexity can lead to challenges in training, risk of overfitting, and may not always result in better accuracy. Due to these pitfalls, backbones with Vision Transformer (ViT) and Convolutional Neural Network (CNN) architectures still outperform best Mamba-based vision models on different vision tasks.
In this work, we systematically re-design the Mamba block to make it more suitable for vision tasks. We propose a hybrid architecture that consists of our proposed formulation (i.e. MambaVision Mixer and MLP) as well as Transformer blocks. Specifically, we study different integration patterns such as adding the Transformer blocks in an iso-parameter manner to earlier, middle and final layers as well as every $l$ layers. Our analysis shows that leveraging several self-attention blocks at the final stages can significantly enhance the capability to capture global context and long-range spatial dependencies. As shown in Sec. 5, using a hybrid architecture also results in a higher image throughput in comparison to both pure Mamba or ViT-based models.
We introduce the MambaVision model which consists of a multi-resolution architecture and leverages CNN-based residual blocks for fast feature extraction of larger resolution features. As shown in Fig. 1, the MambaVision achieves a new SOTA Pareto front in terms of ImageNet-1K Top-1 accuracy and image throughput, outperforming Mamba, CNN and ViT-based models, sometimes by a significant margin. In downstream tasks such as Object detection and instance segmentation as well as semantic segmentation, models with MambaVision backbones outperform comparably-sized counterparts on MS COCO and ADE20 datasets, respectively. Hence, it validates the effectiveness and versatility of MambaVision as an efficient backbone.
To the best of our knowledge, MambaVision is the first effort to study and develop a hybrid architecture comprising of both Mamba and Transformers for computer vision applications. Our main contributions in this work are summarized as follows:
  • We introduce a redesigned vision-friendly Mamba block, improving accuracy and image throughput over the original Mamba architecture.
  • We present a systematic investigation of integration patterns for Mamba and Transformer blocks and demonstrate that incorporating self-attention blocks at the final stages significantly improves the model’s ability to capture global context and long-range spatial dependencies.
  • We introduce MambaVision which is a novel hybrid Mamba Transformer model. The hierarchical MambaVision achieves a new SOTA Pareto front on ImageNet-1K dataset in terms of Top-1 and image throughput tradeoff.
2 Related work
ViT. The Vision Transformer (ViT) [5] emerged as a promising alternative to CNNs, leveraging self-attention layers to provide an enlarged receptive field. However, ViTs initially lacked some of the inherent advantages of CNNs, such as inductive biases and translation invariance, and they required large-scale training datasets to achieve competitive performance. To address these limitations, the Data-efficient Image Transformer (DeiT) [6] introduced a distillation-based training strategy that significantly improved classification accuracy, even with smaller datasets. Building on this, the LeViT [7] model proposed a hybrid approach, incorporating redesigned MLP and self-attention modules optimized for fast inference, enhancing both efficiency and performance. Furthermore, the Cross-covariance Image Transformer (XCiT) [8] introduced a transposed self-attention mechanism, effectively modeling interactions between feature channels and improving the model’s ability to capture complex patterns in the data. The Pyramid Vision Transformer (PVT) [9] adopted a hierarchical structure with patch embedding at the start of each stage and spatial dimension reduction, thereby improving computational efficiency. Similarly, Swin Transformers [10] proposed a layered architecture where self-attention is computed within local windows, which are shifted to enable interaction between regions, balancing local and global contexts. The Twins Transformer [11] featured spatially separable self-attention that significantly enhanced efficiency. In addition, the Focal Transformer [12] utilized focal self-attention for capturing fine-grained details of long-range spatial interactions.
Mamba. Since the introduction of Mamba, a number of efforts have been proposed to leverage its capability for vision applications. Specifically, Vim [3] proposed to use a bidirectional SSM formulation, with the same Mamba formulation, in which tokens were processed in both forward and backward directions in order to capture more global context and improve spatial understanding. However, bidirectional encoding increases the computational load which can slow down training and inference times. In addition, combining information from multiple directions effectively to form a coherent global understanding is challenging as some global context may be lost in the process. As opposed to Vim, our proposed MambaVision uses a single forward pass with a redesigned Mamba block that can capture both short and long-range information and outperform it significantly both in terms of ImageNet Top-1 accuracy and throughput.
EfficientVMamba [4] proposed an atrous-based selective scan with a skip sampling approach to efficiently extract global spatial dependencies. EfficientVMamba also uses a hierarchical architecture, composed of SSM and CNN-based blocks, in which SSMs are used for larger input resolutions to better capture global context while CNNs are used for lower resolutions. In contrast to EfficientVMamba, MambaVision uses CNNs in higher resolutions for faster feature extraction while using both SSMs and self-attention in lower resolution to capture fine-grained details in both short and long-range spatial dependencies. Our proposed MambaVision also outperforms EfficientVMamba by a significant margin in terms of both Top-1 accuracy and image throughput.
In addition, VMamba [13] introduced a generic Mamba-based vision backbone with a Cross-Scan Module (CSM) which facilitates 1D selective scan with an enlarged global receptive field. Specifically, the CSM modules adopt a four-way selective scan methodology (i.e. upper-left and lower-right to opposite directions) to integrate information from all surrounding tokens and capture more global context. In addition, VMamba makes architectural changes such as using depth-wise convolutions and a hierarchical multi-resolution structure. Although the design of the CSM module is more suited for vision tasks, the receptive field is still limited by the cross-scan paths. In comparison to VMamba, the design of our proposed MambaVision mixer is simpler and can capture both short and long-range dependencies. MambaVision also uses CNN-based layer for fast feature extraction as opposed to VMamba which uses the same block structures in all stages. Furthermore, the MambaVision models outperform VMamba counterparts while having significantly higher image throughput.
3 Methdology
3.1 Macro Architecture
In this section, we introduce MambaVision which is our proposed novel architecture with SOTA performance on ImageNet-1K dataset. As illustrated in Fig. 2, MambaVision has a hierarchical architecture consisting of $4$ different stages. The first two stages consist of CNN-based layers for fast feature extraction at higher input resolutions, while stage $3$ and $4$ include the proposed MambaVision and Transformer blocks. Specifically, given an image of size $H\times W\times 3$, the input is first converted into overlapping patches with size $\frac{H}{4}\times\frac{W}{4}\times C$ and projected into a $C$ dimensional embedding space by the stem which consists of two consecutive $3\times 3$ CNN layers with stride of $2$. The downsampler in between stages consists of a batch normalized $3\times 3$ CNN layer with stride $2$ which reduces the image resolution by half. Furthermore, the CNN blocks in stages $1$ and $2$ follow a generic residual block formulation according to the following
Figure 2: The architecture of hierarchical MambaVision models. The first two stages use residual convolutional blocks for fast feature extraction. Stage 3 and 4 employ both MambaVision and Transformer blocks. Specifically, given $N$ layers, we use $\frac{N}{2}$ MambaVision and MLP blocks which are followed by additional $\frac{N}{2}$ Transformer and MLP blocks. The Transformer blocks in final layers allow for recovering lost global context and capture long-range spatial dependencies.
$\begin{split}\mathbf{\hat{z}}&=\text{GELU}(\text{BN}(\text{Conv}_{3\times 3}(\mathbf{z}))),\\ \mathbf{z}&=\text{BN}(\text{Conv}_{3\times 3}(\mathbf{\hat{z}}))+\mathbf{z},\end{split}$ (1)
GELU and BN denote Gaussian Error Linear Unit activation function [14] and batch normalization [15], respectively.
Figure 3: Architecture of MambaVision block. In addition to replacing causal Conv layer with their regular counterparts, we create a symmetric path without SSM as a token mixer to enhance the modeling of global context.
3.2 Micro Architecture
In this section, we first revisit the preliminaries of Mamba and SSMs. We then present the micro design of the architecture in stages 3 and 4 and discuss MambaVision formulation in more details.
3.2.1 Mamba Preliminaries
Mamba is an extension of Structured state space sequence models (S4) which is capable of transforming a 1D continuous input $x(t)\in\mathbb{R}$ to $y(t)\in\mathbb{R}$ via a learnable hidden state $h(t)\in\mathbb{R}^{M}$ with parameters $\bm{A}\in\mathbb{R}^{M\times M}$, $\bm{B}\in\mathbb{R}^{1\times M}$ and $\bm{C}\in\mathbb{R}^{1\times M}$ according to
$\begin{split}h^{\prime}(t)&=\bm{A}h(t)+\bm{B}x(t),\\ y(t)&=\bm{C}h(t),\end{split}$ (2)
Discretization
The continuous parameters $\bm{A}$, $\bm{B}$ and $\bm{C}$ in the above formulation are further converted into discrete parameters for better computational efficiency [16]. Specifically, assuming a timescale $\Delta$, a zero-order hold rule can be applied to obtain discrete parameters $\bm{\bar{A}}\in\mathbb{R}^{M\times M}$, $\bm{\bar{B}}\in\mathbb{R}^{1\times M}$ and $\bm{\bar{C}}\in\mathbb{R}^{1\times M}$ according to
$\bar{\bm{A}}=\exp{(\Delta\bm{A})},$ (3)
$\bar{\bm{B}}=(\Delta\bm{A})^{-1}(\exp{(\Delta\bm{A})}-\bm{I})\cdot(\Delta\bm{B}),$
$\bar{\bm{C}}=\bm{C},$
The Eq. 2 can then be expressed with discrete parameters as
$h(t)=\bar{\bm{A}}h(t-1)+\bar{\bm{B}}x(t),$ (4)
$y(t)=\bar{\bm{C}}h(t),$
In addition, for an input sequence with size $T$, a global convolution with kernel $\bm{\overline{K}}$ can be applied for computing the output of Eq. 4 as in the following
$\begin{split}\bm{\overline{K}}&=(\bm{C}\bm{\overline{B}},\bm{C}\overline{\bm{A}\bm{B}},...,\bm{C}\bm{\overline{A}}^{T-1}\bm{\overline{B}})\\ \quad\bm{y}&=\bm{x}*\bm{\overline{K}},\\ \end{split}$ (5)
Selectivity
Mamba further extends the S4 formulation by introducing a selection mechanism which allows for input-dependant sequence processing. This allows the model’s parameters $\bm{B}$, $\bm{C}$ and $\Delta$ to be adjusted dynamically according to the inputs and filter out irrelevant information. Further discretization details are provided in [2].
3.2.2 Layer Architecture
Assuming an input $X\in\mathbb{R}^{T\times C}$ with sequence length $T$ with embedding dimension $C$, the output of layer $n$ in stages 3 and 4 can be computed as in
$\begin{array}[]{l}\hat{{X}}^{n}=\text{Mixer}(\text{Norm}({X}^{n-1}))+{X}^{n-1}\\ {X}^{n}=\text{MLP}(\text{Norm}(\hat{{X}}^{n}))+\hat{{X}}^{n},\\ \end{array}$ (6)
Norm and Mixer denote the choices of layer normalization and token mixing blocks, respectively. Without loss of generality, Layer Normalization is used for Norm. Given $N$ layers, the first $\frac{N}{2}$ layers employ MambaVision mixer blocks while the remaining $\frac{N}{2}$ layers employ self-attention. We describe the details of each mixer in the following.
Algorithm 1 PyTorch-like pseudo-code for MambaVision mixer
import torch
import math
import torch.nn as nn
import torch.nn.functional as F
from einops import rearrange, repeat
class MambaVisionMixer(nn.Module):
def __init__(self, dim, d_state=16, kernel_size=3):
super().__init__()
self.d_state = d_state
self.dt_rank = math.ceil(dim / 16)
self.in_proj = nn.Linear(self.d_model, self.d_inner)
self.x_proj = nn.Linear(dim//2, self.dt_rank + self.d_state *2)
self.conv1d_x = nn.Conv1d(dim//2, dim//2, kernel_size=kernel_size, padding=’same’, groups=dim//2)
self.conv1d_z = nn.Conv1d(dim//2, dim//2, kernel_size=kernel_size, padding=’same’, groups=dim//2)
self.dt_proj = nn.Linear(self.dt_rank, dim//2)
dt = torch.exp(torch.rand(self.dim//2) * (math.log(dt_max) math.log(dt_min)) + math.log(dt_min)
A_log = torch.log(repeat(torch.arange(1, self.d_state + 1), n -> d n, d=dim//2))
self.A_log = nn.Parameter(A_log)
self.D = nn.Parameter(torch.ones(dim//2))
self.out_proj = nn.Linear(dim, dim)
def forward(self, hidden_states):
xz = rearrange(self.in_proj(hidden_states), b l d -> b d l)
x, z = xz.chunk(2, dim=1)
A = -torch.exp(self.A_log)
x = F.silu(self.conv1d_x(x))
z = F.silu(self.conv1d_z(z))
seqlen = hidden_states.shape[1]
x_dbl = self.x_proj(rearrange(x, b d l -> (b l) d))
dt, B, C = torch.split(x_dbl, [self.dt_rank, self.d_state, self.d_state], dim=-1)
dt = rearrange(self.dt_proj(dt), (b l) d -> b d l, l=seqlen)
B = rearrange(B, (b l) dstate -> b dstate l, l=seqlen)
C = rearrange(C, (b l) dstate -> b dstate l, l=seqlen)
x_ssm = selective_scan_fn(x, dt, A, B, C, D)
hidden_states = rearrange(torch.cat([x_ssm, z], dim=1), b d l -> b l d)
return self.out_proj(hidden_states)
MambaVision Mixer
As shown in Fig. 3, we redesigned the original Mamba mixer to make it more suitable for vision tasks. First, we propose to replace the causal convolution with regular convolution, since it limits the influence to one direction, which is unnecessary and restrictive for vision tasks. In addition, we added a symmetric branch without SSM, consisting of an additional convolution and SiLU activation, to compensate for any content lost due to the sequential constraints of SSMs. We then concatenate the output of both branches and project it via a final linear layer. This combination ensures that the final feature representation incorporates both the sequential and spatial information, leveraging the strengths of both branches. We note that the output of each branch is projected into an embedding space with size $\frac{C}{2}$ (i.e. half the size of original embedding dimension) to maintain similar number of parameters to the original block design. Given an input $X_{in}$, the output of MambaVision mixer $X_{out}$is computed according to
$\begin{array}[]{l}{X_{1}}=\text{Scan}(\sigma(\text{Conv}(\text{Linear}(C,\frac{C}{2})({X_{in}}))))\\ {X_{2}}=\sigma(\text{Conv}(\text{Linear}(C,\frac{C}{2})({X_{in}})))\\ {X_{out}}=\text{Linear}(\frac{C}{2},C)(\text{Concat}(X_{1},X_{2})),\end{array}$ (7)
$\text{Linear}(C_{in},C_{out})(\cdot)$ denotes a linear layer with $C_{in}$ and $C_{out}$ as input and output embedding dimensions, Scan is the selective scan operation as in [2] and $\sigma$ is the activation function for which Sigmoid Linear Unit (SiLU) [17] is used. In addition, Conv and Concat represent 1D convolution and concatenation operations. In Algorithm 1, we present a PyTorch-like pseudo-code for MambaVision mixer. In general, our proposed modification leads to richer feature representations, better generalization, and improved performance on computer vision tasks. We have also experimentally validated the effectiveness of each of our design choices in Sec.
Self-attention
We use a generic multihead self-attention mechanism in accordance to
${\rm Attention}({Q},{K},{V})={\rm Softmax}(\frac{{Q}{K}^{\mathsf{T}}}{\sqrt{d_{h}}}){V}.$ (8)
$Q,K,V$ denote query, key and value respectively and $d_{h}$ is the number of attention heads. Without loss of generality, the attention formulation which can be computed in a windowed manner similar to previous efforts [10, 18]
Table 1: Classification benchmarks on ImageNet-1K [19] dataset. Image throughput is computed on A100 GPU with a batch size of 128.
Model Image Size #Param FLOPs Throughput Top-1
(Px) (M) (G) (Img/Sec) (%)
Conv-Based
ConvNeXt-T [20] 224    28.6   4.5 3196 82.0
ConvNeXt-S [20] 224    50.2   8.7 2008 83.1
ConvNeXt-B [20] 224    88.6 15.4 1485 83.8
RegNetY-040 [21] 288    20.6    6.6 3227 83.0
EfficientNetV2-S [22] 384    21.5    8.0 1735 83.9
Transformer-Based
Swin-T [10] 224    28.3    4.4 2758 81.3
Swin-S [10] 224    49.6    8.5 1720 83.2
Swin-B [10] 224    88.0    15.4 1245 83.5
SwinV2-T [18] 256    28.3    4.4 1674 81.8
SwinV2-S [18] 256    49.7    8.5 1043 83.8
SwinV2-B [18] 256    87.9 15.1    535 84.6
TNT-S [23] 224    23.8    4.8 1478 81.5
Twins-S [11] 224    24.1    2.8 3596 81.7
Twins-B [11] 224    56.1    8.3 1926 83.1
DeiT3-L 224 304.4 59.7   535 84.8
PoolFormer-M58 [24] 224    73.5 11.6    884 82.4
Conv-Transformer
CoaT-Lite-S [25] 224    19.8    4.1 2269 82.3
CrossViT-B [26] 240 105.0 20.1 1321 82.2
NextViT-small 224 31.7 5.8 3834 82.5
NextViT-B 224    44.8   8.3 2926 83.2
MaxViT-B [27] 224    120.0   23.4 507 84.9
MaxViT-L [27] 224    212.0   43.9 376 85.1
FasterViT-1 [28] 224    53.4    5.3 4188 83.2
FasterViT-2 [28] 224    75.9    8.7 3161 84.2
FasterViT-3 [28] 224 159.5 18.2 1780 84.9
Mamba-Based
Vim-T [3] 224 7.0 - 3957 76.1
Vim-S [3] 224 26.0 - 1974 80.5
EfficientVMamba-T [4] 224 6.0 0.8 2904 76.5
EfficientVMamba-S [4] 224 11.0 1.3 1610 78.7
EfficientVMamba-B [4] 224 33.0 4.0 1482 81.8
VMamba-T [13] 224 30.0 4.9 1282 82.6
VMamba-S [13] 224 50.0 8.7 843 83.6
VMamba-B [13] 224 89.0 15.4 645 83.9
MambaVision
MambaVision-T 224   31.8    4.4 6298 82.3
MambaVision-T2 224    35.1    5.1 5990 82.7
MambaVision-S 224    50.1    7.5 4700 83.3
MambaVision-B 224 97.7 15.0 3670 84.2
MambaVision-L 224 227.9 34.9    2190 85.0
MambaVision-L2 224 241.5 37.5    1021 85.3
4 Experiments
Image classification experiments are conducted on ImageNet-1K dataset [19]. We followed the standard training recipe of previous efforts [10, 12, 29] to allow for a comparable analysis for performance of different models. Specifically, all models have been trained for $300$ epochs with a cosine decay scheduler which uses an additional $20$ epochs for warmup and cooldown phases, respectively. Furthermore, we use the LAMB optimizer [30] with global batch size of $4096$, an initial learning rate of $0.005$ and weight decay of $0.05$. We note that using the LAMB optimizer resulted in better results compared to traditional AdamW [31], particularly due to its robustness to higher learning rates. We used 32 A100 GPUs for classification tasks.
To evaluate the performance of downstream tasks, we used our pre-trained models as backbones for object detection and instance segmentation and semantic segmentation tasks and used the MS COCO dataset [32] and ADE20K dataset [33], respectively. Specifically, for object detection and instance segmentation, we used Mask-RCNN [34] head with hyperparameters such as ×$\times 3$ LR schedule with an initial learning rate of $0.0001$, a batch size of $16$ and weight decay of $0.05$. For semantic segmentation, we used an UperNet network [35] head with Adam-W [31] optimizer, an initial learning rate of $56e-$ 5 and a global batch size of 16. We used 8 A100 GPUs for all downstream tasks.
Backbone Param (M) FLOPs (G) $\text{AP}^{\text{box}}$ $\text{AP}^{\text{box}}_{50}$ $\text{AP}^{\text{box}}_{75}$ $\text{AP}^{\text{mask}}$ $\text{AP}^{\text{mask}}_{\text{50}}$ $\text{AP}^{\text{mask}}_{\text{75}}$
Mask-RCNN 3$\times$ schedule
Swin-T [10] 48 267 46.0 68.1 50.3 41.6 65.1 44.9
ConvNeXt-T [20] 48 262 46.2 67.9 50.8 41.7 65.0 44.9
MambaVision-T 48 255 46.4 68.3 51.0 41.8 65.4 45.0
Cascade Mask-RCNN 3$\times$ schedule
DeiT-Small/16 [6] 80 889 48.0 67.2 51.7 41.4 64.2 44.3
ResNet-50 [36] 82 739 46.3 64.3 50.5 40.1 61.7 43.4
Swin-T [10] 86 745 50.4 69.2 54.7 43.7 66.6 47.3
ConvNeXt-T [20] 86 741 50.4 69.1 54.8 43.7 66.5 47.3
MambaVision-T 86 740 51.0 69.9 55.6 44.3 67.2 48.1
X101-32 [37] 101 819 48.1 66.5 52.4 41.6 63.9 45.2
Swin-S [10] 107 838 51.9 70.7 56.3 45.0 68.2 48.8
ConvNeXt-S [20] 108 827 51.9 70.8 56.5 45.0 68.4 49.1
MambaVision-S 108 828 52.1 70.9 56.7 45.2 68.4 49.1
X101-64 [37] 140 972 48.3 66.4 52.3 41.7 64.0 45.1
Swin-B [10] 145 982 51.9 70.5 56.4 45.0 68.1 48.9
ConvNeXt-B [20] 146 964 52.7 71.3 57.2 45.6 68.9 49.5
MambaVision-B 145 964 52.8 71.6 57.2 45.7 69.0 49.5
Table 2: Object detection and instance segmentation results with Mask R-CNN and Cascade Mask R-CNN networks using MS COCO dataset [32].
5 Results
5.1 Image classification
In Table 1, we present the ImageNet-1K classification results. Specifically, we compare against different families of models such as Conv-based, Transformer-based, Conv-Transformer and Mamba-based and demonstrate that our model outperforms the previous efforts by a large margin, considering ImageNet Top-1 accuracy and image throughput. For instance, when compared to popular models such as ConvNeXt and Swin Transformers, MambaVision-B (84.2%) ConvNeXt-B (83.8%) and Swin-B (83.5%), respectively, while also having significantly better image throughput. We observe similar trends in comparison to Mamba-based models. Specifically, MambaVision-B (84.2%) outperforms VMamba-B (83.9%) despite having considerably higher image throughput. We would also like to note that although our main design goal has been to the accuracy and throughput tradeoff, the MambaVision model variants have much lower FLOPs when compared to similarly-sized counterparts. For instance, MambaVision-B has $56\%$ less GFLOPs than MaxViT-B.
5.2 Object Detection and Segmentation
We present object detection and instance segmentation results on MS COCO dataset [32] in Table 2. Specifically, we have trained models with different detection sizes to further validate the effectiveness of MambaVision under different scenarios. We also note that our goal is not to achieve SOTA in these tasks but to compare the performance of our backbone with comparably-sized popular vision backbones and validate its effectiveness. With a simple Mask-RCNN detection head, our pretrained MambaVision-T backbone achieves 46.4 and 41.8 in terms of box AP and mask AP, respectively, outperforming both ConvNeXt-T [20] and Swin-T [10] models. With a Cascade Mask-RCNN network, MambaVision-T, MambaVision-S and MambaVision-B outperform the competing counterparts. Specifically, MambaVision models outperform ConvNeXt-T by +0.6 and 0.6,
Backbone Param (M) FLOPs (G) mIoU
DeiT-Small/16 [6] 52 1099 44.0
Swin-T [10] 60 945 44.5
ResNet-101 [36] 86 1029 44.9
Focal-T [12] 62 998 45.8
Twins-SVT-S [11] 55 - 46.2
MambaVision-T 55 945 46.6
Swin-S [10] 81 1038 47.6
Twins-SVT-B [11] 89 - 47.7
Focal-S [12] 85 1130 48.0
MambaVision-S 84 1135 48.2
Swin-B [10] 121 1188 48.1
Twins-SVT-L [11] 133 - 48.8
Focal-B [12] 126 1354 49.0
MambaVision-B 126 1342 49.1
Table 3: Semantic segmentation results with UperNet [35] networks using ADE20K dataset. For all models, a crop size of $512\times 512$ is used.
ConvNeXt-S by +0.2 and 0.2 and ConvNeXt-B by +0.1 and 0.1 in terms of box AP and mask AP, respectively. Similarly, MambaVision outperforms Swin-T by +0.6 and 0.6, Swin-S by +0.1 and 0.2 and Swin-B by +0.9 and 0.7 in terms of box AP and mask AP, respectively.
We present semantic segmentation benchmarks on the ADE20K dataset [33] in Table 3. For these experiments, we used a UPerNet [35] to be comparable to other models. We observe that MambaVision models outperform similarly-sized competing models for different variants. For instance, MambaVision-T, MambaVision-S and MambaVision-B outperform Swin-T, Swin-S and Swin-B by +0.6, +0.6 and +1.0 in terms of mIoU. Despite the fact that we did not perform extensive optimizations for hyper-parameter tuning of downstream tasks, these results demonstrate the feasibility of MambaVision as a promising backbone for different vision tasks, especially in high-resolution settings.
5.3 Ablation
Design of Token Mixer
In this section, we perform a comprehensive ablation study for a systematic design of MambaVision token mixer. Our goal is to modify the existing Mamba block for computer vision tasks and we evaluate the performance for different tasks such as classification, object detection and instance segmentation and semantic segmentation. All experiments follow a model similar to the architecture layout of the MambaVision-T model as the basis. A shown in Table 4, we start by using the original Mamba formulation with the causal conv layer in the SSM branch (i.e. conv1) and without the additional conv layer in our proposed
  ImageNet COCO ADE20k
[…]
causal conv1 - w/o conv2 80.5 44.8 40.4 44.2
conv1 - w/o conv2 80.9 45.0 40.8 44.7
conv1 - conv2 - w/o concat 81.3 45.3 41.0 45.7
conv1 - conv2 - concat 82.3 46.4 41.8 46.6
Table 4: Systematic design of MambaVision token mixer. w/o and concat refer "without" and concatenation. conv1 and conv2 denote the conv operations in the SSM and additional symmetric branch as shown in Fig. 3. The performance is evaluated on ImageNet-1K, MS COCO and ADE20K dataset for classification, object detection and instance segmentation and semantic segmentation, respectively.
symmetric branch (i.e. conv2). As expected, this formulation achieves a suboptimal performance with a Top-1 accuracy of $80.9\%$ ($-1.8\%$), box AP and mask AP of $44.8$ ($-1.6$) and $40.2$ ($-1.6$) and mIoU of $44.2\%$ ($-1.4$). We then replace the causal conv of SSM branch (i.e. conv1) with a regular
Model Top-1
Random 81.3
First $N/2$ layers (SSSSMMMM) 81.5
Mixed layers-1 (SMSMSMSM) 81.4
Mixed layers-2 (MSMSMSMS) 81.6
Last $N/4$ layers (MMMMMMSS) 81.9
Last $N/2$ layers (MMMMSSSS) 82.3
Table 5: Ablation study of on the effectiveness of different hybrid integration patterns using ImageNet-1K dataset. We use the specified patterns for stages 3 and 4. However, only stage 3 (N=8) patterns have been shown in the table for illustration purposes. S and M denote self-attention and MambaVision token mixer blocks, respectively.
conv layer and observe improvements in terms of all metrics due to this change. Furthermore, we add conv2 layer but use the same gating mechanism of Mamba instead of concatenation. This change improves the performance and leads to a Top-1 accuracy of $81.3\%$, box AP and mask AP of $45.3$ and $41.0$ and mIoU of $45.7\%$. Finally, adding concatenation considerably improves the performance in all metrics by $+1.0\%$, $+1.1$, $+0.8$ and $+0.9$ in terms of ImageNet Top-1, box AP and mask AP for MS COCO and mIoU for ADE20K dataset. Hence, this validates our hypothesis that concatenating the outputs from both branches (i.e. SSM and non-SSM) leads to learning richer feature representation and enhancing global context understanding.
Hybrid Pattern
In this section, we comprehensively study the effect of different hybrid integration patterns for self-attention and MambaVision token mixers. For all experiments, the architecture follows MambaVision-T layout with and we keep the models iso-parameter for fair comparison. The patterns are utilized for stages 3 and 4 with hybrid functionality. We begin our study by using a random pattern and achieve a suboptimal Top-1 accuracy of $81.3\%$. This confirms our previous intuition that simply using self-attention without a specific pattern may not be effective. We then utilize the self-attention blocks in the first $N/2$ layers for each stage, with $N$ representing the total number of stage layers, and observe an improvement of $+0.2\%$ ($81.5\%$) in Top-1 accuracy. However, using a mixed layer pattern of self-attention/MambaVision mixers blocks slightly reduces the accuracy by $-0.1\%$ ($81.4\%$). Conversely, reversing the order of mixed layers by using MambaVision/self-attention improves the performance and archives a Top-1 accuracy of ($81.6\%$). We then utilize self-attention blocks in only the last $N/4$ layers of each stage and observe an immediate improvement of $+0.3\%$ ($81.9\%$) in the accuracy. This validates our hypothesis that using self-attention blocks in the final layers of each stage is an effective design. However, its representation learning capability needs to be tuned with respect to MambaVision layers. Increasing the number of self-attention blocks to $N/2$ last layers of each stage achieves the best performance of $82.3\%$.
6 Conclusion
In this work, we introduced MambaVision which is the first Mamba-Transformer hybrid backbone specifically tailored for vision applications. We proposed re-design of Mamba formulation to enhance global context representation learning capability and presented a comprehensive study of hybrid design integration patterns. MambaVision achieves a new SOTA Pareto front in terms of Top-1 accuracy and image throughput, surpassing Transformer and Mamba-based models by a significant margin. We hope these findings could be the foundation for a new class of hybrid vision models.

### v2

@@
 Abstract
-We propose a novel hybrid Mamba-Transformer backbone, denoted as MambaVision, which is specifically tailored for vision applications.
+We propose a novel hybrid Mamba-Transformer backbone, MambaVision, specifically tailored for vision applications.
 Our core contribution includes redesigning the Mamba formulation to enhance its capability for efficient modeling of visual features.
-In addition, we conduct a comprehensive ablation study on the feasibility of integrating Vision Transformers (ViT) with Mamba.
-Our results demonstrate that equipping the Mamba architecture with several self-attention blocks at the final layers greatly improves the modeling capacity to capture long-range spatial dependencies.
-Based on our findings, we introduce a family of MambaVision models with a hierarchical architecture to meet various design criteria.
-For Image classification on ImageNet-1K dataset, MambaVision model variants achieve a new State-of-the-Art (SOTA) performance in terms of Top-1 accuracy and image throughput.
-In downstream tasks such as object detection, instance segmentation and semantic segmentation on MS COCO and ADE20K datasets, MambaVision outperforms comparably-sized backbones and demonstrates more favorable performance.
-Code: https://github.com/NVlabs/MambaVision.
+Through a comprehensive ablation study, we demonstrate the feasibility of integrating Vision Transformers (ViT) with Mamba.
+Our results show that equipping the Mamba architecture with self-attention blocks in the final layers greatly improves its capacity to capture long-range spatial dependencies.
+Based on these findings, we introduce a family of MambaVision models with a hierarchical architecture to meet various design criteria.
+For classification on the ImageNet-1K dataset, MambaVision variants achieve state-of-the-art (SOTA) performance in terms of both Top-1 accuracy and throughput.
+In downstream tasks such as object detection, instance segmentation, and semantic segmentation on MS COCO and ADE20K datasets, MambaVision outperforms comparably sized backbones while demonstrating favorable performance.
+Code: https://github.com/NVlabs/MambaVision
+[…]
+Figure 1: Top-1 accuracy vs.
+image throughput comparisons on ImageNet-1K dataset.
+The MambaVision models achieve a new Pareto front for Top-1 accuracy and image throughput tradeoff.
+Specifically, MambaVision variants outperform Mamba-based models such as VMamba and Vim, sometimes by a significant margin.
+For all models, image throughput is measured on an A100 NVIDIA GPU with a batch size of 128.
 1 Introduction
-[…]
-Figure 1: Top-1 accuracy vs. image throughput on ImageNet-1K. All measurements conducted on A100 GPU with batch size of 128. MambaVision achieves a new SOTA Pareto front.
-During the recent years, Transformers [1] have become the de facto architecture in different domains including in computer vision, natural language processing, speech processing and robotics.
-In addition, the versatility of the Transformer architecture, primarily due to its attention mechanism, and its flexibility makes it highly suitable for multimodal learning tasks, where integrating and processing information from various modalities is essential.
-Despite these benefits, the quadratic complexity of the attention mechanism with respect to sequence length make Transformers computationally expensive to train and deploy.
+In recent years, Transformers [31] have become the de facto architecture across different domains including computer vision, natural language processing, speech processing, and robotics.
+The versatility of the Transformer architecture, primarily due to its attention mechanism and flexibility, makes it highly suitable for multimodal learning tasks, where integrating and processing information from various modalities is essential.
+Despite these benefits, the quadratic complexity of the attention mechanism with respect to sequence length makes Transformers computationally expensive to train and deploy.
 Recently, Mamba [7] proposed a new State Space Model (SSM) that achieves linear time complexity and outperforms or matches Transformers in different language modeling tasks [7].
 The core contribution of Mamba is a novel selection mechanism that enables efficient input-dependent processing of long sequences with hardware-aware considerations.
-Recently, a number of Mamba-based backbones [3, 4] have also been proposed to leverage the strengths of its SSM formulation in vision tasks such as image classification and semantic segmentation.
+Recently, a number of Mamba-based backbones [40, 25] have been proposed to leverage the strengths of its SSM formulation in vision tasks such as image classification and semantic segmentation.
 However, the Mamba’s autoregressive formulation, while effective for tasks requiring sequential data processing, faces limitations in computer vision tasks that benefit from a full receptive field: (1) Unlike sequences where order matters, image pixels do not have a sequential dependency in the same way.
@@
 MambaVision Mixer and MLP) as well as Transformer blocks.
-Specifically, we study different integration patterns such as adding the Transformer blocks in an iso-parameter manner to earlier, middle and final layers as well as every $l$ layers.
+Specifically, we study different integration patterns such as adding the Transformer blocks in an iso-parameter manner to earlier, middle, and final layers as well as every $l$ layers.
 Our analysis shows that leveraging several self-attention blocks at the final stages can significantly enhance the capability to capture global context and long-range spatial dependencies.
 As shown in Sec.
-5, using a hybrid architecture also results in a higher image throughput in comparison to both pure Mamba or ViT-based models.
+5, using a hybrid architecture also results in higher image throughput compared to both pure Mamba and ViT-based models.
 We introduce the MambaVision model which consists of a multi-resolution architecture and leverages CNN-based residual blocks for fast feature extraction of larger resolution features.
 As shown in Fig.
-1, the MambaVision achieves a new SOTA Pareto front in terms of ImageNet-1K Top-1 accuracy and image throughput, outperforming Mamba, CNN and ViT-based models, sometimes by a significant margin.
-In downstream tasks such as Object detection and instance segmentation as well as semantic segmentation, models with MambaVision backbones outperform comparably-sized counterparts on MS COCO and ADE20 datasets, respectively.
+1, the MambaVision achieves a new SOTA Pareto front in terms of ImageNet-1K Top-1 accuracy and image throughput, outperforming Mamba, CNN, and ViT-based models, sometimes by a significant margin.
+In downstream tasks such as object detection, instance segmentation, and semantic segmentation, models with MambaVision backbones outperform comparably-sized counterparts on MS COCO and ADE20 datasets, respectively.
 Hence, it validates the effectiveness and versatility of MambaVision as an efficient backbone.
@@
   • We introduce a redesigned vision-friendly Mamba block, improving accuracy and image throughput over the original Mamba architecture.
-  • We present a systematic investigation of integration patterns for Mamba and Transformer blocks and demonstrate that incorporating self-attention blocks at the final stages significantly improves the model’s ability to capture global context and long-range spatial dependencies.
-  • We introduce MambaVision which is a novel hybrid Mamba Transformer model.
-The hierarchical MambaVision achieves a new SOTA Pareto front on ImageNet-1K dataset in terms of Top-1 and image throughput tradeoff.
+  • We present a systematic investigation of integration patterns for Mamba and Transformer blocks, and demonstrate that incorporating self-attention blocks at the final stages significantly improves the model’s ability to capture global context and long-range spatial dependencies.
+  • We introduce MambaVision, which is a novel hybrid Mamba-Transformer model.
+The hierarchical MambaVision achieves a new SOTA Pareto front on the ImageNet-1K dataset in terms of Top-1 and image throughput.
+Figure 2: The architecture of hierarchical MambaVision models.
+The first two stages use residual convolutional blocks for fast feature extraction.
+Stages 3 and 4 employ both MambaVision and Transformer blocks.
+Specifically, given $N$ layers, we use $\frac{N}{2}$ MambaVision and MLP blocks, which are followed by additional $\frac{N}{2}$ Transformer and MLP blocks.
+The Transformer blocks in the final layers allow for recovering lost global context and capturing long-range spatial dependencies.
 2 Related work
-ViT.
-The Vision Transformer (ViT) [5] emerged as a promising alternative to CNNs, leveraging self-attention layers to provide an enlarged receptive field.
-However, ViTs initially lacked some of the inherent advantages of CNNs, such as inductive biases and translation invariance, and they required large-scale training datasets to achieve competitive performance.
-To address these limitations, the Data-efficient Image Transformer (DeiT) [6] introduced a distillation-based training strategy that significantly improved classification accuracy, even with smaller datasets.
-Building on this, the LeViT [7] model proposed a hybrid approach, incorporating redesigned MLP and self-attention modules optimized for fast inference, enhancing both efficiency and performance.
-Furthermore, the Cross-covariance Image Transformer (XCiT) [8] introduced a transposed self-attention mechanism, effectively modeling interactions between feature channels and improving the model’s ability to capture complex patterns in the data.
-The Pyramid Vision Transformer (PVT) [9] adopted a hierarchical structure with patch embedding at the start of each stage and spatial dimension reduction, thereby improving computational efficiency.
-Similarly, Swin Transformers [10] proposed a layered architecture where self-attention is computed within local windows, which are shifted to enable interaction between regions, balancing local and global contexts.
-The Twins Transformer [11] featured spatially separable self-attention that significantly enhanced efficiency.
-In addition, the Focal Transformer [12] utilized focal self-attention for capturing fine-grained details of long-range spatial interactions.
-Mamba.
+Conv-Based.
+CNNs have been the cornerstone of computer vision since the introduction of AlexNet [16].
+Recent efforts have focused on modernizing CNN architectures with Transformer-inspired principles.
+ConvNeXt [23] demonstrated competitive performance with Transformers by redesigning ResNet [12] with increased width, larger kernels, and layer normalization.
+RegNetY [26] introduced systematic network design through design space analysis, while EfficientNetV2 [27] leveraged neural architecture search and progressive learning for better efficiency trade-offs.
+Despite strong performance, these CNN models inherently lack the global receptive field needed for capturing long-range dependencies.
+Transformer-Based.
+ViTs [5] marked a significant shift in computer vision by introducing self-attention mechanisms for enlarged receptive fields.
+However, ViTs initially lacked the inherent advantages of CNNs and required extensive training data.
+To address these limitations, DeiT [28] introduced distillation-based training for improved accuracy on smaller datasets, while Swin Transformer [21] proposed a hierarchical architecture using shifted windows for self-attention, effectively balancing local and global context.
+Models such as Twins [3] and PVT [32] further enhanced efficiency through spatially separable self-attention and hierarchical structures with patch embedding.
+Despite these advances, the quadratic complexity of self-attention operations in these models continued to pose efficiency challenges.
+Conv-Transformer.
+The complementary strengths of CNNs and ViTs inspired hybrid architectures.
+CoAT [36] and CrossViT [1] demonstrated enhanced feature learning by combining convolutions with self-attention, while NextViT [17] systematically incorporated CNN-like processing into Transformers.
+Recent efforts like EfficientFormer [18] and FasterViT [10] focused on optimizing efficiency-accuracy trade-offs, achieving competitive performance with high throughput through carefully designed hybrid architectures.
+Mamba-Based.
 Since the introduction of Mamba, a number of efforts have been proposed to leverage its capability for vision applications.
-Specifically, Vim [3] proposed to use a bidirectional SSM formulation, with the same Mamba formulation, in which tokens were processed in both forward and backward directions in order to capture more global context and improve spatial understanding.
-However, bidirectional encoding increases the computational load which can slow down training and inference times.
-In addition, combining information from multiple directions effectively to form a coherent global understanding is challenging as some global context may be lost in the process.
-As opposed to Vim, our proposed MambaVision uses a single forward pass with a redesigned Mamba block that can capture both short and long-range information and outperform it significantly both in terms of ImageNet Top-1 accuracy and throughput.
-EfficientVMamba [4] proposed an atrous-based selective scan with a skip sampling approach to efficiently extract global spatial dependencies.
-EfficientVMamba also uses a hierarchical architecture, composed of SSM and CNN-based blocks, in which SSMs are used for larger input resolutions to better capture global context while CNNs are used for lower resolutions.
-In contrast to EfficientVMamba, MambaVision uses CNNs in higher resolutions for faster feature extraction while using both SSMs and self-attention in lower resolution to capture fine-grained details in both short and long-range spatial dependencies.
-Our proposed MambaVision also outperforms EfficientVMamba by a significant margin in terms of both Top-1 accuracy and image throughput.
-In addition, VMamba [13] introduced a generic Mamba-based vision backbone with a Cross-Scan Module (CSM) which facilitates 1D selective scan with an enlarged global receptive field.
-Specifically, the CSM modules adopt a four-way selective scan methodology (i.e.
-upper-left and lower-right to opposite directions) to integrate information from all surrounding tokens and capture more global context.
-In addition, VMamba makes architectural changes such as using depth-wise convolutions and a hierarchical multi-resolution structure.
-Although the design of the CSM module is more suited for vision tasks, the receptive field is still limited by the cross-scan paths.
-In comparison to VMamba, the design of our proposed MambaVision mixer is simpler and can capture both short and long-range dependencies.
-MambaVision also uses CNN-based layer for fast feature extraction as opposed to VMamba which uses the same block structures in all stages.
-Furthermore, the MambaVision models outperform VMamba counterparts while having significantly higher image throughput.
-3 Methdology
+Vim [40] introduced a bidirectional SSM formulation that processes tokens in both forward and backward directions to capture global context and improve spatial understanding.
+However, this bidirectional approach faces significant limitations: it increases computational overhead, slows down training and inference times, and struggles to effectively combine information from multiple directions without losing global context.
+In contrast, MambaVision achieves superior results using a single forward pass with our redesigned Mamba block, demonstrating significantly better ImageNet Top-1 accuracy and throughput.
+VMamba [20] proposed a generic Mamba-based vision backbone featuring a Cross-Scan Module (CSM).
+This module implements a four-way selective scan methodology (from upper-left and lower-right to opposite directions) to integrate information from surrounding tokens and capture global context.
+VMamba also incorporates architectural modifications like depth-wise convolutions and a hierarchical multi-resolution structure.
+While the CSM module is designed for vision tasks, its receptive field remains constrained by the cross-scan paths.
+MambaVision offers several advantages over VMamba: our mixer design is simpler yet captures both short and long-range dependencies more effectively, we employ CNN-based layers for fast feature extraction rather than using uniform block structures across all stages, and we achieve superior performance with significantly higher throughput.
+EfficientVMamba [25] uses SSMs for larger resolutions and CNNs for lower ones, while MambaVision takes the opposite approach with CNNs at higher resolutions and SSM/self-attention at lower ones, leading to significantly better accuracy and throughput.
+Similarly, while SiMBA [24] addresses Mamba’s stability through EinFFT channel modeling, it doesn’t fully address spatial understanding limitations, resulting in lower performance compared to MambaVision’s comprehensive design.
+3 Methodology
 3.1 Macro Architecture
@@
 Specifically, given an image of size $H\times W\times 3$, the input is first converted into overlapping patches with size $\frac{H}{4}\times\frac{W}{4}\times C$ and projected into a $C$ dimensional embedding space by the stem which consists of two consecutive $3\times 3$ CNN layers with stride of $2$.
-The downsampler in between stages consists of a batch normalized $3\times 3$ CNN layer with stride $2$ which reduces the image resolution by half.
+The downsampler in between stages consists of a $3\times 3$ CNN layer with stride $2$ which reduces the image resolution by half.
 Furthermore, the CNN blocks in stages $1$ and $2$ follow a generic residual block formulation according to the following
-Figure 2: The architecture of hierarchical MambaVision models.
-The first two stages use residual convolutional blocks for fast feature extraction.
-Stage 3 and 4 employ both MambaVision and Transformer blocks.
-Specifically, given $N$ layers, we use $\frac{N}{2}$ MambaVision and MLP blocks which are followed by additional $\frac{N}{2}$ Transformer and MLP blocks.
-The Transformer blocks in final layers allow for recovering lost global context and capture long-range spatial dependencies.
 $\begin{split}\mathbf{\hat{z}}&=\text{GELU}(\text{BN}(\text{Conv}_{3\times 3}(\mathbf{z}))),\\ \mathbf{z}&=\text{BN}(\text{Conv}_{3\times 3}(\mathbf{\hat{z}}))+\mathbf{z},\end{split}$ (1)
 GELU and BN denote Gaussian Error Linear Unit activation function [14] and batch normalization [15], respectively.
-Figure 3: Architecture of MambaVision block.
-In addition to replacing causal Conv layer with their regular counterparts, we create a symmetric path without SSM as a token mixer to enhance the modeling of global context.
-3.2 Micro Architecture
-In this section, we first revisit the preliminaries of Mamba and SSMs. We then present the micro design of the architecture in stages 3 and 4 and discuss MambaVision formulation in more details.
-3.2.1 Mamba Preliminaries
-Mamba is an extension of Structured state space sequence models (S4) which is capable of transforming a 1D continuous input $x(t)\in\mathbb{R}$ to $y(t)\in\mathbb{R}$ via a learnable hidden state $h(t)\in\mathbb{R}^{M}$ with parameters $\bm{A}\in\mathbb{R}^{M\times M}$, $\bm{B}\in\mathbb{R}^{1\times M}$ and $\bm{C}\in\mathbb{R}^{1\times M}$ according to
-$\begin{split}h^{\prime}(t)&=\bm{A}h(t)+\bm{B}x(t),\\ y(t)&=\bm{C}h(t),\end{split}$ (2)
-Discretization
-The continuous parameters $\bm{A}$, $\bm{B}$ and $\bm{C}$ in the above formulation are further converted into discrete parameters for better computational efficiency [16].
-Specifically, assuming a timescale $\Delta$, a zero-order hold rule can be applied to obtain discrete parameters $\bm{\bar{A}}\in\mathbb{R}^{M\times M}$, $\bm{\bar{B}}\in\mathbb{R}^{1\times M}$ and $\bm{\bar{C}}\in\mathbb{R}^{1\times M}$ according to
-$\bar{\bm{A}}=\exp{(\Delta\bm{A})},$ (3)
-$\bar{\bm{B}}=(\Delta\bm{A})^{-1}(\exp{(\Delta\bm{A})}-\bm{I})\cdot(\Delta\bm{B}),$
-$\bar{\bm{C}}=\bm{C},$
-The Eq. 2 can then be expressed with discrete parameters as
-$h(t)=\bar{\bm{A}}h(t-1)+\bar{\bm{B}}x(t),$ (4)
-$y(t)=\bar{\bm{C}}h(t),$
-In addition, for an input sequence with size $T$, a global convolution with kernel $\bm{\overline{K}}$ can be applied for computing the output of Eq. 4 as in the following
-$\begin{split}\bm{\overline{K}}&=(\bm{C}\bm{\overline{B}},\bm{C}\overline{\bm{A}\bm{B}},...,\bm{C}\bm{\overline{A}}^{T-1}\bm{\overline{B}})\\ \quad\bm{y}&=\bm{x}*\bm{\overline{K}},\\ \end{split}$ (5)
-Selectivity
-Mamba further extends the S4 formulation by introducing a selection mechanism which allows for input-dependant sequence processing.
-This allows the model’s parameters $\bm{B}$, $\bm{C}$ and $\Delta$ to be adjusted dynamically according to the inputs and filter out irrelevant information.
-Further discretization details are provided in [2].
-3.2.2 Layer Architecture
-Assuming an input $X\in\mathbb{R}^{T\times C}$ with sequence length $T$ with embedding dimension $C$, the output of layer $n$ in stages 3 and 4 can be computed as in
-$\begin{array}[]{l}\hat{{X}}^{n}=\text{Mixer}(\text{Norm}({X}^{n-1}))+{X}^{n-1}\\ {X}^{n}=\text{MLP}(\text{Norm}(\hat{{X}}^{n}))+\hat{{X}}^{n},\\ \end{array}$ (6)
-Norm and Mixer denote the choices of layer normalization and token mixing blocks, respectively.
-Without loss of generality, Layer Normalization is used for Norm.
-Given $N$ layers, the first $\frac{N}{2}$ layers employ MambaVision mixer blocks while the remaining $\frac{N}{2}$ layers employ self-attention.
-We describe the details of each mixer in the following.
+Please see the supplementary materials for further details regarding MambaVision macro architecture.
 Algorithm 1 PyTorch-like pseudo-code for MambaVision mixer
@@
 self.dt_proj = nn.Linear(self.dt_rank, dim//2)
-dt = torch.exp(torch.rand(self.dim//2) * (math.log(dt_max) math.log(dt_min)) + math.log(dt_min)
+dt = torch.exp(torch.rand(self.dim//2) * (math.log(dt_max) - math.log(dt_min)) + math.log(dt_min))
 A_log = torch.log(repeat(torch.arange(1, self.d_state + 1), n -> d n, d=dim//2))
@@
 return self.out_proj(hidden_states)
+3.2 Micro Architecture
+In this section, we first revisit the preliminaries of Mamba and SSMs. We then present the micro design of the architecture in stages 3 and 4 and discuss MambaVision formulation in more details.
+3.2.1 Mamba Preliminaries
+In Mamba, a 1D continuous input $x(t)\in\mathbb{R}$ is transformed into $y(t)\in\mathbb{R}$ via a learnable hidden state $h(t)\in\mathbb{R}^{M}$ with parameters $\bm{A}\in\mathbb{R}^{M\times M}$, $\bm{B}\in\mathbb{R}^{M\times 1}$ and $\bm{C}\in\mathbb{R}^{1\times M}$ according to
+$\begin{split}h^{\prime}(t)&=\bm{A}h(t)+\bm{B}x(t),\\ y(t)&=\bm{C}h(t),\end{split}$ (2)
+Discretization
+The continuous parameters $\bm{A}$, $\bm{B}$ and $\bm{C}$ in the above formulation are further converted into discrete parameters for better computational efficiency [8].
+Specifically, assuming a timescale $\Delta$, a zero-order hold rule can be applied to obtain discrete parameters $\bm{\bar{A}}\in\mathbb{R}^{M\times M}$, $\bm{\bar{B}}\in\mathbb{R}^{M\times 1}$ and $\bm{\bar{C}}\in\mathbb{R}^{1\times M}$ according to
+$\bar{\bm{A}}=\exp{(\Delta\bm{A})},$ (3)
+$\bar{\bm{B}}=(\Delta\bm{A})^{-1}(\exp{(\Delta\bm{A})}-\bm{I})\cdot(\Delta\bm{B}),$
+$\bar{\bm{C}}=\bm{C},$
+The Eq. 2 can then be expressed with discrete parameters as
+$h(t)=\bar{\bm{A}}h(t-1)+\bar{\bm{B}}x(t),$ (4)
+$y(t)=\bar{\bm{C}}h(t),$
+In addition, for an input sequence with size $T$, a global convolution with kernel $\bm{\overline{K}}$ can be applied for computing the output of Eq. 4 as in the following
+$\begin{split}\bm{\overline{K}}&=(\bm{C}\bm{\overline{B}},\bm{C}\overline{\bm{A}\bm{B}},...,\bm{C}\bm{\overline{A}}^{T-1}\bm{\overline{B}}),\\ \quad\bm{y}&=\bm{x}*\bm{\overline{K}},\\ \end{split}$ (5)
+Selectivity
+Mamba further extends the S4 formulation by introducing a selection mechanism which allows for input-dependant sequence processing.
+This allows the model’s parameters $\bm{B}$, $\bm{C}$ and $\Delta$ to be adjusted dynamically according to the inputs and filter out irrelevant information.
+Further discretization details are provided in [7].
+3.2.2 Layer Architecture
+Assuming an input $X\in\mathbb{R}^{T\times C}$ with sequence length $T$ with embedding dimension $C$, the output of layer $n$ in stages 3 and 4 can be computed as in
+$\begin{array}[]{l}\hat{{X}}^{n}=\text{Mixer}(\text{Norm}({X}^{n-1}))+{X}^{n-1},\\ {X}^{n}=\text{MLP}(\text{Norm}(\hat{{X}}^{n}))+\hat{{X}}^{n},\\ \end{array}$ (6)
+Norm and Mixer denote the choices of layer normalization and token mixing blocks, respectively.
+Without loss of generality, Layer Normalization is used for Norm.
+Given $N$ layers, the first $\frac{N}{2}$ layers employ MambaVision mixer blocks while the remaining $\frac{N}{2}$ layers employ self-attention.
+We describe the details of each mixer in the following.
+Figure 3: Architecture of MambaVision block.
+In addition to replacing causal Conv layer with their regular counterparts, we create a symmetric path without SSM as a token mixer to enhance the modeling of global context.
 MambaVision Mixer
@@
 First, we propose to replace the causal convolution with regular convolution, since it limits the influence to one direction, which is unnecessary and restrictive for vision tasks.
-In addition, we added a symmetric branch without SSM, consisting of an additional convolution and SiLU activation, to compensate for any content lost due to the sequential constraints of SSMs.
+In addition, we added a symmetric branch without SSM, consisting of an additional convolution and Sigmoid Linear Unit (SiLU) [6] activation, to compensate for any content lost due to the sequential constraints of SSMs.
 We then concatenate the output of both branches and project it via a final linear layer.
@@
 Given an input $X_{in}$, the output of MambaVision mixer $X_{out}$ is computed according to
-$\begin{array}[]{l}{X_{1}}=\text{Scan}(\sigma(\text{Conv}(\text{Linear}(C,\frac{C}{2})({X_{in}}))))\\ {X_{2}}=\sigma(\text{Conv}(\text{Linear}(C,\frac{C}{2})({X_{in}})))\\ {X_{out}}=\text{Linear}(\frac{C}{2},C)(\text{Concat}(X_{1},X_{2})),\end{array}$ (7)
-$\text{Linear}(C_{in},C_{out})(\cdot)$ denotes a linear layer with $C_{in}$ and $C_{out}$ as input and output embedding dimensions, Scan is the selective scan operation as in [2] and $\sigma$ is the activation function for which Sigmoid Linear Unit (SiLU) [17] is used.
+$\begin{array}[]{l}{X_{1}}=\text{Scan}(\sigma(\text{Conv}(\text{Linear}(C,\frac{C}{2})({X_{in}})))),\\ {X_{2}}=\sigma(\text{Conv}(\text{Linear}(C,\frac{C}{2})({X_{in}}))),\\ {X_{out}}=\text{Linear}(\frac{C}{2},C)(\text{Concat}(X_{1},X_{2})),\end{array}$ (7)
+$\text{Linear}(C_{in},C_{out})(\cdot)$ denotes a linear layer with $C_{in}$ and $C_{out}$ as input and output embedding dimensions, Scan is the selective scan operation as in [7] and $\sigma$ is the activation function for which SiLU is used.
 In addition, Conv and Concat represent 1D convolution and concatenation operations.
@@
 We have also experimentally validated the effectiveness of each of our design choices in Sec.
+5.3.
 Self-attention
@@
 $Q,K,V$ denote query, key and value respectively and $d_{h}$ is the number of attention heads.
-Without loss of generality, the attention formulation which can be computed in a windowed manner similar to previous efforts [10, 18]
-Table 1: Classification benchmarks on ImageNet-1K [19] dataset. Image throughput is computed on A100 GPU with a batch size of 128.
-Model Image Size #Param FLOPs Throughput Top-1
+In addition, our framework allows for computing the attention in a windowed manner similar to previous efforts [21, 22] (see the supplementary materials for window size ablation study).
+4 Experiments
+Image classification experiments are conducted on the ImageNet-1K dataset [4].
+We followed the standard training recipe of previous efforts [21, 37, 11] to allow for a comparable analysis of performance across different models.
+Specifically, all models have been trained for 300 epochs using 32 A100 GPUs.
+The self-attention formulation in stages 3 and 4 of all MambaVision variants use a window size of 14 and 7, respectively.
+For detailed training configurations, please see the provided anonymous code repository.
+To evaluate the performance of downstream tasks, we used our pre-trained models as backbones for object detection, instance segmentation, and semantic segmentation tasks using the MS COCO dataset [19] and ADE20K dataset [39], respectively.
+Specifically, for object detection and instance segmentation, we used Cascade Mask-RCNN [13] head with hyperparameters such as ×3 LR schedule.
+For semantic segmentation, we used a UperNet network [34] head and 8 A100 GPUs for all experiments.
+Table 1: Comparison of classification benchmarks on ImageNet-1K dataset [4]. Image throughput is measured on A100 GPU with a batch size of 128.
+Model Image Size #Params FLOPs Throughput Top-1
 (Px) (M) (G) (Img/Sec) (%)
@@
 RegNetY-040 [26] 288    20.6    6.6 3227 83.0
+ResNetV2-101 [33] 224    44.5    7.8 4019 82.0
 EfficientNetV2-S [27] 384    21.5    8.0 1735 83.9
@@
 Swin-S [21] 224    49.6    8.5 1720 83.2
-Swin-B [10] 224    88.0    15.4 1245 83.5
 SwinV2-T [22] 256    28.3    4.4 1674 81.8
@@
 Twins-B [3] 224    56.1    8.3 1926 83.1
-DeiT3-L 224 304.4 59.7   535 84.8
+Twins-L [3] 224    99.3 14.8 1439 83.7
+DeiT-B [28] 224    86.6 16.9 2035 82.0
+DeiT3-L [29] 224 304.4 59.7   535 84.8
 PoolFormer-M58 [38] 224    73.5 11.6    884 82.4
@@
 CoaT-Lite-S [36] 224    19.8    4.1 2269 82.3
+CrossViT-S [1] 240 26.9 5.1 2832 81.0
 CrossViT-B [1] 240 105.0 20.1 1321 82.2
-NextViT-small 224 31.7 5.8 3834 82.5
-NextViT-B 224    44.8   8.3 2926 83.2
+Visformer-S [2] 224    40.2    4.8 3676 82.1
+NextViT-S [17] 224 31.7 5.8 3834 82.5
+NextViT-B [17] 224    44.8   8.3 2926 83.2
+NextViT-L [17] 224    57.8 10.8 2360 83.6
+EfficientFormer-L1 [18] 224 12.3 1.31 6220 79.2
+EfficientFormer-L3 [18] 224    31.4   3.9 2845 82.4
+EfficientFormer-L7 [18] 224    82.2 10.2 1359 83.4
 MaxViT-B [30] 224    120.0   23.4 507 84.9
@@
 EfficientVMamba-B [25] 224 33.0 4.0 1482 81.8
+SiMBA-S [24] 224 15.3 2.4 826 81.7
+SiMBA-B [24] 224 22.8 4.2 624 83.5
 VMamba-T [20] 224 30.0 4.9 1282 82.6
@@
 MambaVision-L2 224 241.5 37.5    1021 85.3
-4 Experiments
-Image classification experiments are conducted on ImageNet-1K dataset [19].
-We followed the standard training recipe of previous efforts [10, 12, 29] to allow for a comparable analysis for performance of different models.
-Specifically, all models have been trained for $300$ epochs with a cosine decay scheduler which uses an additional $20$ epochs for warmup and cooldown phases, respectively.
-Furthermore, we use the LAMB optimizer [30] with global batch size of $4096$, an initial learning rate of $0.005$ and weight decay of $0.05$.
-We note that using the LAMB optimizer resulted in better results compared to traditional AdamW [31], particularly due to its robustness to higher learning rates.
-We used 32 A100 GPUs for classification tasks.
-To evaluate the performance of downstream tasks, we used our pre-trained models as backbones for object detection and instance segmentation and semantic segmentation tasks and used the MS COCO dataset [32] and ADE20K dataset [33], respectively.
-Specifically, for object detection and instance segmentation, we used Mask-RCNN [34] head with hyperparameters such as ×$\times 3$ LR schedule with an initial learning rate of $0.0001$, a batch size of $16$ and weight decay of $0.05$.
-For semantic segmentation, we used an UperNet network [35] head with Adam-W [31] optimizer, an initial learning rate of $56e-$ 5 and a global batch size of 16.
-We used 8 A100 GPUs for all downstream tasks.
-Backbone Param (M) FLOPs (G) $\text{AP}^{\text{box}}$ $\text{AP}^{\text{box}}_{50}$ $\text{AP}^{\text{box}}_{75}$ $\text{AP}^{\text{mask}}$ $\text{AP}^{\text{mask}}_{\text{50}}$ $\text{AP}^{\text{mask}}_{\text{75}}$
-Mask-RCNN 3$\times$ schedule
-Swin-T [10] 48 267 46.0 68.1 50.3 41.6 65.1 44.9
-ConvNeXt-T [20] 48 262 46.2 67.9 50.8 41.7 65.0 44.9
-MambaVision-T 48 255 46.4 68.3 51.0 41.8 65.4 45.0
-Cascade Mask-RCNN 3$\times$ schedule
+Backbone Params (M) FLOPs (G) $\text{AP}^{\text{box}}$ $\text{AP}^{\text{box}}_{50}$ $\text{AP}^{\text{box}}_{75}$ $\text{AP}^{\text{mask}}$ $\text{AP}^{\text{mask}}_{\text{50}}$ $\text{AP}^{\text{mask}}_{\text{75}}$
 DeiT-Small/16 [28] 80 889 48.0 67.2 51.7 41.4 64.2 44.3
@@
 ConvNeXt-T [23] 86 741 50.4 69.1 54.8 43.7 66.5 47.3
-MambaVision-T 86 740 51.0 69.9 55.6 44.3 67.2 48.1
+MambaVision-T 86 740 51.1 70.0 55.6 44.3 67.3 47.9
 X101-32 [35] 101 819 48.1 66.5 52.4 41.6 63.9 45.2
@@
 ConvNeXt-S [23] 108 827 51.9 70.8 56.5 45.0 68.4 49.1
-MambaVision-S 108 828 52.1 70.9 56.7 45.2 68.4 49.1
+MambaVision-S 108 828 52.3 71.1 56.7 45.2 68.5 48.9
 X101-64 [35] 140 972 48.3 66.4 52.3 41.7 64.0 45.1
@@
 ConvNeXt-B [23] 146 964 52.7 71.3 57.2 45.6 68.9 49.5
-MambaVision-B 145 964 52.8 71.6 57.2 45.7 69.0 49.5
-Table 2: Object detection and instance segmentation results with Mask R-CNN and Cascade Mask R-CNN networks using MS COCO dataset [32].
+MambaVision-B 145 964 52.8 71.3 57.2 45.7 68.7 49.4
+Table 2: Object detection and instance segmentation benchmarks using Cascade Mask R-CNN [13] on MS COCO dataset [19].
+All models are trained by using a 3×3$\times$3 × schedule and a crop resolution of $1280\times 800$.
 5 Results
@@
 In Table 1, we present the ImageNet-1K classification results.
-Specifically, we compare against different families of models such as Conv-based, Transformer-based, Conv-Transformer and Mamba-based and demonstrate that our model outperforms the previous efforts by a large margin, considering ImageNet Top-1 accuracy and image throughput.
-For instance, when compared to popular models such as ConvNeXt and Swin Transformers, MambaVision-B (84.2%) ConvNeXt-B (83.8%) and Swin-B (83.5%), respectively, while also having significantly better image throughput.
+Specifically, we compare against different families of models such as Conv-based, Transformer-based, Conv-Transformer, and Mamba-based architectures and demonstrate that our model outperforms the previous efforts by a large margin, considering ImageNet Top-1 accuracy and image throughput.
+For example, MambaVision-B achieves higher accuracy (84.2%) compared to ConvNeXt-B (83.8%) and Swin-B (83.5%), while also having significantly better image throughput.
 We observe similar trends in comparison to Mamba-based models.
 Specifically, MambaVision-B (84.2%) outperforms VMamba-B (83.9%) despite having considerably higher image throughput.
-We would also like to note that although our main design goal has been to the accuracy and throughput tradeoff, the MambaVision model variants have much lower FLOPs when compared to similarly-sized counterparts.
-For instance, MambaVision-B has $56\%$ less GFLOPs than MaxViT-B.
+We observe similar trends in performance comparisons with respect to other Mamba-based models.
+In addition, we would also like to note that although our main design goal has been to optimize the accuracy and throughput tradeoff, the MambaVision model variants have much lower FLOPs when compared to similarly-sized counterparts.
+For instance, MambaVision-B has 56% less GFLOPs than MaxViT-B.
 5.2 Object Detection and Segmentation
-We present object detection and instance segmentation results on MS COCO dataset [32] in Table 2.
-Specifically, we have trained models with different detection sizes to further validate the effectiveness of MambaVision under different scenarios.
-We also note that our goal is not to achieve SOTA in these tasks but to compare the performance of our backbone with comparably-sized popular vision backbones and validate its effectiveness.
-With a simple Mask-RCNN detection head, our pretrained MambaVision-T backbone achieves 46.4 and 41.8 in terms of box AP and mask AP, respectively, outperforming both ConvNeXt-T [20] and Swin-T [10] models.
-With a Cascade Mask-RCNN network, MambaVision-T, MambaVision-S and MambaVision-B outperform the competing counterparts.
-Specifically, MambaVision models outperform ConvNeXt-T by +0.6 and 0.6,
+We evaluate our model’s object detection and instance segmentation performance on the MS COCO dataset [19], as shown in Table 2.
+To comprehensively validate MambaVision’s effectiveness, we trained models of varying sizes and compared them with popular vision backbones of comparable scale under identical conditions.
+Using a Cascade Mask R-CNN [13] head, all variants of MambaVision demonstrated superior performance compared to their counterparts.
+Specifically, MambaVision models outperform ConvNeXt-T by +0.7 and +0.6, ConvNeXt-S by +0.4 and +0.2 and ConvNeXt-B by +0.1 and +0.1 in terms of box Average Precision (AP) and mask AP, respectively.
+Similarly, MambaVision outperforms Swin-T by +0.7 and +0.6, Swin-S by +0.4 and +0.2 and Swin-B by +0.9 and +0.7 in terms of box AP and mask AP, respectively.
+For semantic segmentation, we evaluated the performance on the ADE20K dataset [39] using UPerNet [34], as shown in Table 3.
+We observe that MambaVision models outperform similarly-sized competing models for different variants.
+For instance, MambaVision-T, MambaVision-S, and MambaVision-B outperform Swin-T, Swin-S, and Swin-B by +1.5, +0.6, and +1.0 in terms of mIoU, respectively.
+Notably, these improvements were achieved without extensive hyperparameter optimization for downstream tasks, highlighting MambaVision’s potential as a robust backbone for various vision tasks, particularly in high-resolution scenarios.
+Moreover, our approach consistently attains higher mIoU than Focal Transformers across all scales while having comparable model sizes.
 Backbone Param (M) FLOPs (G) mIoU
@@
 Focal-T [37] 62 998 45.8
-Twins-SVT-S [11] 55 - 46.2
-MambaVision-T 55 945 46.6
+MambaVision-T 55 945 46.0
 Swin-S [21] 81 1038 47.6
@@
 MambaVision-B 126 1342 49.1
-Table 3: Semantic segmentation results with UperNet [35] networks using ADE20K dataset. For all models, a crop size of $512\times 512$ is used.
-ConvNeXt-S by +0.2 and 0.2 and ConvNeXt-B by +0.1 and 0.1 in terms of box AP and mask AP, respectively.
-Similarly, MambaVision outperforms Swin-T by +0.6 and 0.6, Swin-S by +0.1 and 0.2 and Swin-B by +0.9 and 0.7 in terms of box AP and mask AP, respectively.
-We present semantic segmentation benchmarks on the ADE20K dataset [33] in Table 3.
-For these experiments, we used a UPerNet [35] to be comparable to other models.
-We observe that MambaVision models outperform similarly-sized competing models for different variants.
-For instance, MambaVision-T, MambaVision-S and MambaVision-B outperform Swin-T, Swin-S and Swin-B by +0.6, +0.6 and +1.0 in terms of mIoU.
-Despite the fact that we did not perform extensive optimizations for hyper-parameter tuning of downstream tasks, these results demonstrate the feasibility of MambaVision as a promising backbone for different vision tasks, especially in high-resolution settings.
+Table 3: Semantic segmentation results with UperNet [34] model using ADE20K dataset. All models are trained using a crop resolution of $512\times 512$.
 5.3 Ablation
+Large-scale Training on ImageNet-21K
+For the first time in any Mamba-based approach, our work (MambaVision) has scaled training to the large ImageNet-21K dataset with significantly bigger model sizes.
+As demonstrated in Fig.
+4 the results are promising.
+Specifically, we observe meaningful improvements for the smaller MambaVision-B model (97.7M parameters), whose Top-1 accuracy increases from 84.2% to 84.9% at 224 resolution.
+In addition, pre-training and fine-tuning MambaVision-L raises its Top-1 accuracy from 85% to 86.1% at 224 resolution.
+We have also introduced a larger variant, MambaVision-L3 (739.6M parameters), which attains Top-1 accuracies of 87.3% and 88.1% at 256 and 512 resolutions, respectively.
+These results validate the scalability of our model across larger datasets, varying model sizes, and different image resolutions.
+To the best of our knowledge, MambaVision represents the first successful scaling of a Mamba-based vision architecture to ImageNet-21K with strong performance.
+This ability to scale is critical for real-world scenarios that rely on massive datasets, where bigger, more capable models are needed to achieve robust performance.
+We anticipate that MambaVision’s proven scalability will further encourage the adoption of Mamba-based models in industrial and large-scale research applications.
 Design of Token Mixer
-In this section, we perform a comprehensive ablation study for a systematic design of MambaVision token mixer.
-Our goal is to modify the existing Mamba block for computer vision tasks and we evaluate the performance for different tasks such as classification, object detection and instance segmentation and semantic segmentation.
-All experiments follow a model similar to the architecture layout of the MambaVision-T model as the basis.
-A shown in Table 4, we start by using the original Mamba formulation with the causal conv layer in the SSM branch (i.e.
-conv1) and without the additional conv layer in our proposed
+We conducted a comprehensive ablation study to systematically design the MambaVision token mixer.
+Our investigation focused on adapting the Mamba block for computer vision tasks, evaluating performance across classification, object detection, instance segmentation, and semantic segmentation.
+All experiments used a model architecture based on MambaVision-T configuration.
+As shown in Table 4, we began with the original Mamba formulation, which includes a causal convolution layer in the SSM branch (conv1) but lacks the additional convolution layer in our proposed symmetric branch (conv2).
+This baseline configuration achieved suboptimal performance across all metrics, with ImageNet Top-1 accuracy of 80.9% (-1.8%), MS COCO box AP of 44.8 (-1.6) and mask AP of 40.2 (-1.6), and ADE20K mIoU of 44.2% (-1.4).
+We then replaced the causal convolution in the SSM branch (conv1) with a regular convolution layer, which improved performance across all metrics.
+Subsequently, we added conv2 layer while maintaining Mamba’s original gating mechanism instead of concatenation, resulting in ImageNet Top-1 accuracy of 81.3%, MS COCO box AP of 45.3 and mask AP of 41.0, and ADE20K mIoU of 45.7%.
+Finally, implementing concatenation led to substantial improvements across all metrics, with gains of +1.0% in ImageNet Top-1, +1.1 in box AP and +0.8 in mask AP for MS COCO, and +0.9 in mIoU for ADE20K.
+These results validate our hypothesis that concatenating outputs from both branches (SSM and non-SSM) enables the model to learn richer feature representations and enhance global context understanding.
+Figure 4: Performance scalability of MambaVision ImageNet-21K pretrained models with varying model sizes and resolutions.
+Hybrid Pattern
+We conducted a comprehensive study examining various hybrid integration patterns between self-attention and MambaVision token mixers.
+All experiments maintained the MambaVision-T architecture layout with iso-parameter models for fair comparison, implementing hybrid functionality in stages 3 and 4.
+Initial experiments with a random integration pattern yielded suboptimal results with a Top-1 accuracy of 81.3%, confirming our intuition that arbitrary self-attention placement may be ineffective.
+When we positioned self-attention blocks in the first N/2 layers of each stage (where N represents the total number of stage layers), performance improved by +0.2% (81.5%).
+A mixed layer pattern alternating between self-attention and MambaVision mixer blocks showed a slight performance decrease of -0.1% (81.4%), while reversing this order to MambaVision/self-attention improved accuracy to 81.6%.
+Placing self-attention blocks in only the last N/4 layers of each stage yielded a significant improvement of +0.3% (81.9%), supporting our hypothesis that self-attention is most effective in the final layers.
+Further optimization revealed that extending self-attention to the last N/2 layers of each stage achieved the best performance at 82.3%, indicating the importance of carefully balancing self-attention blocks with MambaVision layers for optimal representation learning.
+Table 4: Systematic design of MambaVision token mixer.
+w/o and concat refer to "without" and concatenation.
+Conv1 and conv2 denote the conv operations in the SSM and additional symmetric branch as shown in Fig.
+3.
+COCO experiments are performed using Mask-RCNN [13] head and ×1 LR schedule.
   ImageNet COCO ADE20k
@@
 conv1 - conv2 - w/o concat 81.3 45.3 41.0 45.7
-conv1 - conv2 - concat 82.3 46.4 41.8 46.6
-Table 4: Systematic design of MambaVision token mixer.
-w/o and concat refer "without" and concatenation.
-conv1 and conv2 denote the conv operations in the SSM and additional symmetric branch as shown in Fig.
-3.
-The performance is evaluated on ImageNet-1K, MS COCO and ADE20K dataset for classification, object detection and instance segmentation and semantic segmentation, respectively.
-symmetric branch (i.e.
-conv2).
-As expected, this formulation achieves a suboptimal performance with a Top-1 accuracy of $80.9\%$ ($-1.8\%$), box AP and mask AP of $44.8$ ($-1.6$) and $40.2$ ($-1.6$) and mIoU of $44.2\%$ ($-1.4$).
-We then replace the causal conv of SSM branch (i.e.
-conv1) with a regular
-Model Top-1
-Random 81.3
-First $N/2$ layers (SSSSMMMM) 81.5
-Mixed layers-1 (SMSMSMSM) 81.4
-Mixed layers-2 (MSMSMSMS) 81.6
-Last $N/4$ layers (MMMMMMSS) 81.9
-Last $N/2$ layers (MMMMSSSS) 82.3
-Table 5: Ablation study of on the effectiveness of different hybrid integration patterns using ImageNet-1K dataset.
-We use the specified patterns for stages 3 and 4.
-However, only stage 3 (N=8) patterns have been shown in the table for illustration purposes.
-S and M denote self-attention and MambaVision token mixer blocks, respectively.
-conv layer and observe improvements in terms of all metrics due to this change.
-Furthermore, we add conv2 layer but use the same gating mechanism of Mamba instead of concatenation.
-This change improves the performance and leads to a Top-1 accuracy of $81.3\%$, box AP and mask AP of $45.3$ and $41.0$ and mIoU of $45.7\%$.
-Finally, adding concatenation considerably improves the performance in all metrics by $+1.0\%$, $+1.1$, $+0.8$ and $+0.9$ in terms of ImageNet Top-1, box AP and mask AP for MS COCO and mIoU for ADE20K dataset.
-Hence, this validates our hypothesis that concatenating the outputs from both branches (i.e.
-SSM and non-SSM) leads to learning richer feature representation and enhancing global context understanding.
-Hybrid Pattern
-In this section, we comprehensively study the effect of different hybrid integration patterns for self-attention and MambaVision token mixers.
-For all experiments, the architecture follows MambaVision-T layout with and we keep the models iso-parameter for fair comparison.
-The patterns are utilized for stages 3 and 4 with hybrid functionality.
-We begin our study by using a random pattern and achieve a suboptimal Top-1 accuracy of $81.3\%$.
-This confirms our previous intuition that simply using self-attention without a specific pattern may not be effective.
-We then utilize the self-attention blocks in the first $N/2$ layers for each stage, with $N$ representing the total number of stage layers, and observe an improvement of $+0.2\%$ ($81.5\%$) in Top-1 accuracy.
-However, using a mixed layer pattern of self-attention/MambaVision mixers blocks slightly reduces the accuracy by $-0.1\%$ ($81.4\%$).
-Conversely, reversing the order of mixed layers by using MambaVision/self-attention improves the performance and archives a Top-1 accuracy of ($81.6\%$).
-We then utilize self-attention blocks in only the last $N/4$ layers of each stage and observe an immediate improvement of $+0.3\%$ ($81.9\%$) in the accuracy.
-This validates our hypothesis that using self-attention blocks in the final layers of each stage is an effective design.
-However, its representation learning capability needs to be tuned with respect to MambaVision layers.
-Increasing the number of self-attention blocks to $N/2$ last layers of each stage achieves the best performance of $82.3\%$.
+conv1 - conv2 - concat 82.3 46.4 41.8 46.0
+Figure 5: Visualizations of MambaVision’s self-attention layers showing how the model learns to focus on semantically meaningful regions via attention maps (middle) and overlays (right).
+Table 5: Ablation study of on the effectiveness of different hybrid integration patterns. S and M denote self-attention and MambaVision token mixer blocks, respectively.
+Model Pattern Params (M) Top-1
+Random - 31.8 81.3
+First $N/2$ layers SSSSMMMM 31.8 81.5
+Mixed layers-1 SMSMSMSM 31.8 81.4
+Mixed layers-2 MSMSMSMS 31.8 81.6
+Last $N/4$ layers MMMMMMSS 31.8 81.9
+Last $N/2$ layers MMMMSSSS 31.8 82.3
+Interpretability
+To better understand how MambaVision processes visual information, we visualize the attention maps from the self-attention layers in the final stages.
+As shown in Fig.
+5, these visualizations reveal that the model learns to focus on semantically meaningful regions without explicit supervision.
+In the aircraft example, the attention clearly highlights the entire plane body, suggesting effective capture of object boundaries.
+For the bird image, we observe concentrated attention on distinctive features like the head and tail regions, demonstrating the model’s ability to identify fine-grained details.
+In the case of object-human interaction (bottom row), the attention map shows strong activation on both the subject and the object being held, indicating that the self-attention layers successfully model relationships between different elements in the scene.
+These visualizations support our architectural design choice of using self-attention blocks in the final stages to capture global context and long-range dependencies.
 6 Conclusion
 In this work, we introduced MambaVision which is the first Mamba-Transformer hybrid backbone specifically tailored for vision applications.
-We proposed re-design of Mamba formulation to enhance global context representation learning capability and presented a comprehensive study of hybrid design integration patterns.
-MambaVision achieves a new SOTA Pareto front in terms of Top-1 accuracy and image throughput, surpassing Transformer and Mamba-based models by a significant margin.
+We proposed re-design of Mamba formulation to enhance global context representation learning capability.
+MambaVision achieves a new SOTA Pareto front in terms of Top-1 accuracy and image throughput, outperforming Transformer and Mamba-based models by a significant margin.
+Through extensive experimentation across multiple vision tasks, including classification, detection and segmentation, we demonstrated the versatility and effectiveness of our approach.
+Our systematic analysis of integration patterns revealed that positioning self-attention blocks in the final layers significantly improves the model’s ability to capture long-range dependencies while maintaining efficiency.
+Furthermore, we successfully scaled MambaVision to ImageNet-21K pretraining, achieving strong performance that matches SOTA models, demonstrating its potential for large-scale vision applications.
+The success of MambaVision in addressing the limitations of pure Mamba-based architectures while leveraging their strengths opens new possibilities for vision backbone design.
 We hope these findings could be the foundation for a new class of hybrid vision models.
+Appendix
+Figure S.1: Visualization of MambaVision’s attention patterns.
+Each row contains two example cases, with each case showing a triplet of: (left) original input image, (middle) attention heat map, and (right) attention overlay on the input image.
+The examples showcase diverse scenarios: containers and spiders (row 1), aircraft and birds (row 2), marine life and snakes (row 3), groomed dogs and extreme sports (row 4), poultry and snakes (row 5), and arachnids and outdoor activities (row 6).
+The attention maps reveal how MambaVision effectively localizes key semantic regions and object boundaries across this wide range of categories.
+G Ablation Study
+To determine the optimal window size for MambaVision models, we study its impact on the performance of MambaVision-T in different tasks such as image classification, object detection and instance segmentation.
+Given $Q,K,V$ as the query, key and value tensors respectively, self-attention is computed according to
+${\rm Attention}({Q},{K},{V})={\rm Softmax}(\frac{{Q}{K}^{\mathsf{T}}}{\sqrt{d_{h}}}){V}.$ (9)
+$d_{h}$ represents the number of attention heads.
+If the input size is larger than the window size, the attention is computed in the local windows.
+Specifically, we study two different architectures with window sizes of 7 and 14 in their stage 3 of the model.
+We also measure image throughput for the task of image classification with a batch size of 128.
+As presented in Table S.1, our analysis reveals that increasing the window size to 14 offers a favorable trade-off between performance and computational cost.
+While maintaining nearly identical throughput (6298 img/s vs.
+6318 img/s), the larger window size achieves consistent improvements across all vision benchmarks: ImageNet top-1 accuracy increases to 82.3%, COCO mask AP improves to 41.8%.
+These gains, though modest, come with minimal computational overhead on modern hardware such as the NVIDIA A100 GPU.
+Based on this empirical evidence, we selected 14 and 7 as our default window sizes, as this combination provides better vision understanding capabilities while preserving the model’s efficiency.
+The negligible 0.3% decrease in throughput is well justified by the improved performance in various vision tasks.
+  Model Window Size Throughput (img/s) ImageNet COCO
+[…]
+MambaVision-T 7,7 6318 82.2 46.4 41.7
+MambaVision-T 14,7 6298 82.3 46.4 41.8
+Table S.1: Ablation study on window size for MambaVision model’s performance.
+Experiments on COCO dataset [19] are performed using Mask-RCNN [13] head and ×1 LR schedule.
+Throughput is measured for image classification on a single NVIDIA A100 GPU with batch size 128.
+H Architecture Details
+In Table S.2, we present the comprehensive architectural specifications of MambaVision variants.
+The backbone follows a hierarchical design with 4 stages, each employing convolutional down-sampling operations that progressively reduce spatial resolution by a factor of two.
+A key innovation in our architecture appears in Stages 3 and 4, where we introduce a hybrid design that synergistically combines Mamba-based sequence modeling with self-attention mechanisms.
+This hybrid approach leverages Mamba’s efficient sequence processing capabilities while benefiting from the global context modeling strengths of self-attention layers.
+Each variant (T, S, B, and L) maintains this fundamental structure while scaling the channel dimensions and layer counts to achieve different complexity-performance trade-offs.
+Output Size
+(Downs. Rate)
+MambaVision-T MambaVision-S MambaVision-B MambaVision-L
+Stem 112$\times$112 (2$\times$) $\begin{bmatrix}\text{Conv-BN-ReLU}\\ \text{C:32, S:2}\end{bmatrix}$$\times$ 1 $\begin{bmatrix}\text{Conv-BN-ReLU}\\ \text{C:64, S:2}\end{bmatrix}$$\times$ 1 $\begin{bmatrix}\text{Conv-BN-ReLU}\\ \text{C:64, S:2}\end{bmatrix}$$\times$ 1 $\begin{bmatrix}\text{Conv-BN-ReLU}\\ \text{C:64, S:2}\end{bmatrix}$$\times$ 1
+$\begin{bmatrix}\text{Conv-BN-ReLU}\\ \text{C:80}\end{bmatrix}$$\times$ 1 $\begin{bmatrix}\text{Conv-BN-ReLU}\\ \text{C:96}\end{bmatrix}$$\times$ 1 $\begin{bmatrix}\text{Conv-BN-ReLU}\\ \text{C:128}\end{bmatrix}$$\times$ 1 $\begin{bmatrix}\text{Conv-BN-ReLU}\\ \text{C:196}\end{bmatrix}$$\times$ 1
+Stage 1 56$\times$56 (4$\times$) Conv, C:160, S:2 Conv, C:192, S:2 Conv, C:256, S:2 Conv, C:392, S:2
+$\begin{bmatrix}\text{ResBlock}\\ \text{C:160}\end{bmatrix}$$\times$1, $\begin{bmatrix}\text{ResBlock}\\ \text{C:192}\end{bmatrix}$$\times$ 3, $\begin{bmatrix}\text{ResBlock}\\ \text{C:256}\end{bmatrix}$$\times$ 3, $\begin{bmatrix}\text{ResBlock}\\ \text{C:392}\end{bmatrix}$$\times$ 3,
+Stage 2 28$\times$28 (8$\times$) Conv, C:320, S:2 Conv, C:384, S:2 Conv, C:512, S:2 Conv, C:768, S:2
+$\begin{bmatrix}\text{ResBlock}\\ \text{C:320}\end{bmatrix}$$\times$ 3, $\begin{bmatrix}\text{ResBlock}\\ \text{C:384}\end{bmatrix}$$\times$ 3, $\begin{bmatrix}\text{ResBlock}\\ \text{C:512}\end{bmatrix}$$\times$ 3, $\begin{bmatrix}\text{ResBlock}\\ \text{C:768}\end{bmatrix}$$\times$ 3,
+Stage 3 14$\times$14 (16$\times$) Conv, C:640, S:2 Conv, C:768, S:2 Conv, C:1024, S:2 Conv, C:1568, S:2
+$\begin{bmatrix}\text{MV}\\ \text{C:640}\end{bmatrix}$$\times$ 4, $\begin{bmatrix}\text{SA}\\ \text{C:640, head:8}\end{bmatrix}$$\times$ 4 , $\begin{bmatrix}\text{MV}\\ \text{C:768}\end{bmatrix}$$\times$ 4,$\begin{bmatrix}\text{SA}\\ \text{C:768, head:8}\end{bmatrix}$$\times$ 3, $\begin{bmatrix}\text{MV}\\ \text{C:1024}\end{bmatrix}$$\times$ 4,$\begin{bmatrix}\text{SA}\\ \text{C:1024, head:8}\end{bmatrix}$$\times$ 4, $\begin{bmatrix}\text{MV}\\ \text{C:1568}\end{bmatrix}$$\times$ 4,$\begin{bmatrix}\text{SA}\\ \text{C:1568, head:16}\end{bmatrix}$$\times$ 3,
+Stage 4 7$\times$7 (32$\times$) Conv, C:1280, S:2 Conv, C:1536, S:2 Conv, C:2048, S:2 Conv, C:3136, S:2
+$\begin{bmatrix}\text{MV}\\ \text{C:1280}\end{bmatrix}$$\times$ 4,$\begin{bmatrix}\text{SA}\\ \text{C:1280, head:16}\end{bmatrix}$$\times$ 4, $\begin{bmatrix}\text{MV}\\ \text{C:1536}\end{bmatrix}$$\times$ 4,$\begin{bmatrix}\text{SA}\\ \text{C:1536, head:16}\end{bmatrix}$$\times$ 2, $\begin{bmatrix}\text{MV}\\ \text{C:2048}\end{bmatrix}$$\times$ 4,$\begin{bmatrix}\text{SA}\\ \text{C:2048, head:16}\end{bmatrix}$$\times$ 2, $\begin{bmatrix}\text{MV}\\ \text{C:3136}\end{bmatrix}$$\times$ 4,$\begin{bmatrix}\text{SA}\\ \text{C:3136, head:32}\end{bmatrix}$$\times$ 2,
+Table S.2: Architecture configurations of MambaVision models. SA and MV refer to self-attention and MambaVision mixer blocks respectively. BN denote Batch Normalization.
+I Training Details
+Image classification experiments are conducted on the ImageNet-1K dataset [4].
+All models have been trained for 300 epochs using 32 A100 GPUs, with LAMB optimizer, batch size of 4096, and learning rate of 4e-3.
+The self-attention formulation in stages 3 and 4 of all MambaVision variants use a window size of 14 and 7, respectively.
+To evaluate the performance of downstream tasks, we used our pre-trained models as backbones for object detection, instance segmentation, and semantic segmentation tasks using the MS COCO dataset [19] and ADE20K dataset [39], respectively.
+For all downstream tasks, we used an AdamW optimizer and batch size of 16.
+Specifically, for object detection and instance segmentation, we used the Cascade Mask-RCNN [13] head with hyperparameters such as ×3 LR schedule.
+For semantic segmentation, we used a UperNet network [34] segmentation head.
+J Interpretability
+To demonstrate the interpretability of MambaVision models, we visualize the attention patterns learned by our model across diverse object categories.
+Figure S.1 presents a comprehensive analysis of attention mechanisms through paired examples.
+Our paired visualization analysis reveals several key insights about MambaVision’s visual processing capabilities:
+  • Consistent Pattern Recognition: Each triplet (input-heatmap-overlay) demonstrates how the model maintains consistent attention patterns across different instances of similar object categories.
+  • Contextual Understanding: The paired examples within each row often represent contrasting scenarios (e.g., man-made objects vs. natural subjects), showing the model’s adaptability across domains.
+  • Fine-grained Detail: The attention heat maps precisely highlight discriminative features, from the texture of animal fur to the structural elements of vehicles and containers.
+  • Robust Localization: Across all example pairs, the overlaid visualizations demonstrate accurate object boundary detection, regardless of the subject’s position or background complexity.